# Topologically Driven Spin-Orbit Torque in Dirac Matter

### v1

@@ Abstract
We unveil novel spin-orbit torque mechanisms driven by topological edge states in magnetic graphene-based devices. Within the energy gap, a damping-like torque plateau emerges within the quantum anomalous Hall phase upon breaking particle-hole symmetry, while for energies at the spin-split Dirac points located within the bands, a large damping-like torque develops as a result of a vanishing Fermi contour. Such torques are tunable by the degree of spin-pseudospin entanglement dictated by proximity-induced spin-orbit coupling terms. This additionally allows to reach the upper limit of charge-to-spin conversion in non-magnetic devices.
The spin-orbit torque (SOT) mechanism represents a novel method to electrically manipulate a magnetic state, providing power-efficiency and miniaturization prospects beyond traditional multi-ferromagnet setups [1]. While the traditional spin-transfer torque mechanism relies on the generation of spin-polarized currents using ferromagnets [2], the SOT mechanism leverages SOC within metals to generate a current-driven spin source via charge-to-spin conversion [3, 4]. The non-equilibrium spin source $\boldsymbol{S}$ couples to the magnetic state, oriented along $\hat{\boldsymbol{m}}$, exerting a torque $\boldsymbol{T}\propto\hat{\boldsymbol{m}}\times\boldsymbol{S}$, driving the magnetization dynamics to eventually flip the magnetic state. The conventional SOT phenomenon focuses on heavy metal/ferromagnet bilayers, considering two contributions: the damping-like (DL) torque which stems from a spin current injected from the bulk metal to the ferromagnet via the spin Hall effect, and the field-like (FL) torque generated by an interfacial spin density via the Rashba-Edelstein effect (REE).
@@
Maximally efficient charge-to-spin conversion.— We first demonstrate the impact of spin-pseudospin entanglement on charge-to-spin conversion by analyzing the REE. Here, broken mirror symmetry parallel to the 2D plane generates helical spin-momentum locking, which under the current-driven carrier redistribution gives rise to an in-plane spin density perpendicular to the applied current, of the form $S_{\text{FL}}\propto\hat{\boldsymbol{z}}\times\boldsymbol{\mathcal{E}}$, with $\boldsymbol{\mathcal{E}}$ the driving electric field. [24]. Vertical Rashba splitting in graphene generates favorable conditions for the REE, as transport occurs through a single spin-helical band [22]. However, because pseudospin mediates spin-momentum locking the spin texture is quenched by spin-pseudospin entanglement [23], representing the main limiting factor in this system. Manipulating spin-pseudospin entanglement thus emerges as a path towards optimizing charge-to-spin conversion, which we now demonstrate through the interplay between Rashba SOC and Kane-Mele SOC.
We consider a minimal graphene model comprising proximity-induced Rashba SOC and Kane-Mele SOC. The Hamiltonian is given by
$\begin{split}\mathcal{H}_{\boldsymbol{k}}=&\,\hbar v\left(\tau k_{x}\sigma_{x}+k_{y}\sigma_{y}\right)-\frac{\lambda_{\text{R}}}{2}\left(s_{x}\sigma_{y}-\tau s_{y}\sigma_{x}\right)\\ &-\frac{\lambda_{\text{KM}}}{2}\tau s_{z}\sigma_{z}\text{ ,}\end{split}$ (1)
@@
At large momentum the bands are completely pseudospin polarized along the radial direction ±$\pm\hat{\boldsymbol{k}}$, with each subset of bands splitting into clockwise and counter-clockwise spin helicities along ±$\pm\hat{\boldsymbol{\varphi}}=\pm\hat{\boldsymbol{z}}\times\hat{\boldsymbol{k}}$, resulting in orthogonal pure states of the form $|\mu\hat{\boldsymbol{\varphi}},\nu\hat{\boldsymbol{k}}\rangle$, where the first and second components respectively correspond to the spin and pseudospin substates, with $\mu,\,\nu=\pm$. However, near the Dirac point Rashba SOC mixes states of opposite polarizations, promoting two set of entangled states of the form $(|\mu\hat{\boldsymbol{\varphi}},\nu\hat{\boldsymbol{k}}\rangle\pm|-\mu\hat{\boldsymbol{\varphi}},-\nu\hat{\boldsymbol{k}}\rangle)/\sqrt{2}$. At this point spin-pseudospin entanglement acts as a double-edged sword as it enables transport across a single spin-helical band by opening a pseudogap, at the expense of quenching the spin texture. Kane-Mele SOC allows us to wield this sword to our advantage as it induces opposite effects on each set of bands. Indeed, in one set of bands the pseudogap magnitude increases along with the entanglement, while in the remaining bands entanglement decreases thus enhancing the spin helicity throughout the pseudogap [25]. This effect is depicted in the left panel of Fig. 1.
Figure 1: Left panel: Band structure for non-magnetic graphene with $\lambda_{\text{KM}}=\lambda_{\text{R}}=4\,\text{meV}$, with maximal spin helicity $s_{\varphi}$ throughout the entire Rashba pseudogap (yellow highlight). Right panel: REE efficiency $\Theta_{\text{REE}}$ showing maximal efficiency for $\lambda_{\text{KM}}=\lambda_{\text{R}}$ (solid curve).
@@
$\boldsymbol{S}_{\text{surf}}(\varepsilon_{\text{F}})=\sum_{n}\int\frac{\text{d}^{2}\boldsymbol{k}}{(2\pi)^{2}}\,\delta\mathit{f}_{n,\boldsymbol{k}}\boldsymbol{s}_{n,\boldsymbol{k}}\text{ ,}$ (2)
where $\varepsilon_{\text{F}}$ is the Fermi level, $\boldsymbol{s}_{n,\boldsymbol{k}}\equiv\langle E_{n,\boldsymbol{k}}|\boldsymbol{s}|E_{n,\boldsymbol{k}}\rangle$ is the spin texture, with $n$ the band index and $|E_{n,\boldsymbol{k}}\rangle$ the state of band $n$ at momentum $\boldsymbol{k}$, and $\delta\mathit{f}_{n,\boldsymbol{k}}$ is the current-induced variation of the carrier occupation. The REE efficiency, $\Theta_{\text{REE}}$, is defined as the adimensional ratio between the non-equilibrium spin density $S_{\text{FL}}$ and the longitudinal conductivity $\sigma$, $\Theta_{\text{REE}}=(2v_{\text{F}}S_{\text{FL}})/(\hbar\sigma)$, which is limited to $|\Theta_{\text{REE}}|\leq 1$, with $v_{\text{F}}$ the Fermi velocity [22]. Within the pseudogap $\Theta_{\text{REE}}$ is solely determined by the spin helicity, thus we conclude that the REE is optimized by adequately manipulating the spin-pseudospin entanglement of the mid-pseudogap Bloch states. This is achieved by tuning Rashba SOC and Kane-Mele SOC, which enhances the spin helicity of the band that crosses the pseudogap, while additionally opening a wider spectral range for optimal charge-to-spin conversion (see supplementary material in Ref. [25]). A prior study predicted a large $\Theta_{\text{REE}}$ for a purely Rashba graphene system arguing that Kane-Mele SOC is negligible in graphene/transition metal dichalcogenide bilayers [22]. Here we show that Kane-Mele SOC allows to surpass previous proposals, accessing the upper limit of charge-to-spin conversion efficiency, where the REE efficiency reaches its maximal value of $|\Theta_{\text{REE}}|=1$ for $|\lambda_{\text{KM}}|=|\lambda_{\text{R}}|$ (see right panel of Fig.1), indicating that all carriers are fully spin polarized.
Topologically driven spin-orbit torque.— In SOT devices the spin polarization is coupled to a magnetic state, which is captured in the minimal model Hamiltonian by including the term $-\frac{1}{2}J_{\text{ex}}\hat{\boldsymbol{m}}\cdot\boldsymbol{s}$, representing an exchange splitting $J_{\text{ex}}$ along the magnetization direction $\hat{\boldsymbol{m}}$, which we take along the out-of-plane axis $\hat{\boldsymbol{z}}$. Such an effect is obtained by proximitizing graphene with an insulating ferromagnet [37, 38, 39, 40, 41, 42]. The exchange energy is usually the dominant energy scale, splitting the energy dispersion into spin majority and spin minority Dirac cones centered about energy $\mp J_{\text{ex}}/2$, while Rashba SOC opens a gap at the crossing between both cones and induces a helical component to the spin texture, as shown in Fig. 2-(a).
When including a magnetization, in addition to the FL torque generated by the REE, a DL torque of the form $S_{\text{DL}}\propto m_{z}\boldsymbol{\mathcal{E}}$ emerges from the entire Fermi-sea contribution of non-equilibrium spin textures. The Fermi sea non-equilibrium spin density is calculated as
$\boldsymbol{S}_{\text{sea}}(\varepsilon_{\text{F}})=\sum_{n}\int\frac{\text{d}^{2}\boldsymbol{k}}{(2\pi)^{2}}\,\mathit{f}_{n,\boldsymbol{k}}\delta\boldsymbol{s}_{n,\boldsymbol{k}}\text{ ,}$ (3)
where $\mathit{f}_{n,\boldsymbol{k}}$ is the equilibrium carrier occupation, and $\delta\boldsymbol{s}_{n,\boldsymbol{k}}$ corresponds to a non-equilibrium component of the spin texture. We first describe this mechanism in the simple case of a 2D electron gas, as proposed in Ref. [13]. When applying an electric field, $\boldsymbol{\mathcal{E}}$, the wave vector of the Bloch electrons is modified as $\hbar\frac{\text{d}}{\text{d}t}\boldsymbol{k}=-e\boldsymbol{\mathcal{E}}$ (with $e$ the elementary charge), thus generating a non-equilibrium effective magnetic field stemming from spin-momentum locking in a moving rest-frame. The interaction between the equilibrium spin polarization and the non-equilibrium field yields a non-equilibrium component to the spin texture throughout the entire Fermi sea, captured by $\delta\boldsymbol{s}_{\boldsymbol{k}}$ [13].
Figure 2: (a) Band structure exhibiting spin-split Dirac cones at energies $\mp J_{\text{ex}}/2$ along the out-of-plane axis $\hat{\boldsymbol{z}}$, and a band gap at charge neutrality (yellow highlight), with $J_{\text{ex}}=40\,\text{meV}$ and $\lambda_{\text{KM}}=\lambda_{\text{R}}=4\,\text{meV}$ (b) Spin texture of the valence band, showing a sharp quenching at the band edge (yellow highlight) due to Kane-Mele SOC. (c) Damping-like non-equilibrium spin texture (arbitrary units) of the valence band, showing a source of DL torque at the band edge and at the Dirac point.
@@
We offer an explanation for the origin of the SOT plateau based on the semi-classical spin-pseudospin dynamics, relating it to band structure features. As shown in Fig. 2-(a), the band gap at charge neutrality (yellow highlight) is generated by the avoided crossing between the upper band of the spin majority Dirac cone and the lower band of the spin minority one, with positive and negative out-of-plane spin polarization respectively and both exhibiting the same helicity. For $\lambda_{\text{KM}}=0$ these bands present a radial pseudospin polarization along ±$\pm\hat{\boldsymbol{k}}$, opposite to one another, belonging to orthogonal pseudospin subspaces. Thus, the hybridization between both bands near the gap leads to a loss of their out-of-plane spin polarization due to spin-pseudospin entanglement, while their in-plane spin texture is unaffected. On the other hand, when introducing Kane-Mele SOC both bands acquire the same out-of-plane pseudospin component proportional to $\lambda_{\text{KM}}$, and thus share a common pseudospin subspace. In this case the hybridization between bands of up and down spin polarization stemming from the opposite Dirac cones modifies the in-plane spin texture [25]. This process results in a sharp variation of the spin helicity at the band edges, as highlighted in yellow in Fig. 2-(b), where the mark of strong spin-momentum locking forecasts a large $S_{\text{DL}}$. Indeed, as the Bloch states are transported under an applied electric field, those near the band edge feel a strong variation of the effective magnetic field, which drives the non-equilibrium spin dynamics resulting in a non-equilibrium spin texture source at the band edge, shown in Fig. 2-(c). This effect is opposite at the valence and conduction bands, manifesting broken particle-hole symmetry, and therefore, within the gap the fully filled valence band generates the DL SOT plateau [25].

### v2

@@ Abstract
We unveil novel spin-orbit torque mechanisms driven by topological edge states in magnetic graphene-based devices. Within the energy gap, a damping-like torque plateau emerges within the quantum anomalous Hall phase upon breaking particle-hole symmetry, while for energies at the spin-split Dirac points located within the bands, a large damping-like torque develops as a result of a vanishing Fermi contour. Such torques are tunable by the degree of spin-pseudospin entanglement dictated by proximity-induced spin-orbit coupling terms.
The spin-orbit torque (SOT) mechanism represents a novel method to electrically manipulate a magnetic state, providing power-efficiency and miniaturization prospects beyond traditional multi-ferromagnet setups [1]. While the traditional spin-transfer torque mechanism relies on the generation of spin-polarized currents using ferromagnets [2], the SOT mechanism leverages SOC within metals to generate a current-driven spin source via charge-to-spin conversion [3, 4]. The non-equilibrium spin source $\bm{S}$ couples to the magnetic state, oriented along $\hat{\bm{m}}$, exerting a torque $\bm{T}\propto\hat{\bm{m}}\times\bm{S}$, driving the magnetization dynamics to eventually flip the magnetic state. The conventional SOT phenomenon focuses on heavy metal/ferromagnet bilayers, considering two contributions: the damping-like (DL) torque which stems from a spin current injected from the bulk metal to the ferromagnet via the spin Hall effect, and the field-like (FL) torque generated by an interfacial spin density via the Rashba-Edelstein effect (REE).
@@
We consider a minimal graphene model comprising both Rashba and Kane-Mele SOC with an out-of-plane magnetization. The Hamiltonian is given by
@@
with $v\sim 10^{6}\,\text{m/s}$ the velocity of massless Dirac electrons in graphene, $\lambda_{\text{R}}$ and $\lambda_{\text{KM}}$ the Rashba and Kane-Mele SOC parameters respectively, and $J_{\text{ex}}$ the exchange splitting along the magnetization direction $\hat{\bm{z}}$. The spin and pseudospin degrees of freedom are respectively represented by the Pauli vectors $\bm{s}$ and $\bm{\sigma}$, while $\tau=\pm$ represents the valley index, where we will focus our analysis on the $\tau=+$ subspace without loss of generality [24]. While SOC in bare graphene is negligible, such SOC fields can be induced by proximity effects in graphene/topological insulator [25, 26, 27, 28, 29] and graphene/transition metal dichalcogenide heterostructures [30, 31, 32, 33, 34, 35]. On the other hand, the magnetization term is obtained by interfacing graphene with an insulating ferromagnet [36, 37, 38, 39, 40, 41].
The exchange splitting is usually the dominant energy scale, splitting the energy dispersion into spin majority and spin minority Dirac cones centered about energy $\mp J_{\text{ex}}/2$. Rashba SOC opens a gap at the crossing between both cones and induces a helical component to the spin texture, as shown in Fig. 1-(a).
@@
$\bm{S}_{\text{surf}}(\varepsilon_{\text{F}})=\sum_{n}\int\frac{\text{d}^{2}\bm{k}}{(2\pi)^{2}}\,\delta\mathit{f}_{n,\bm{k}}\bm{s}_{n,\bm{k}}\text{ ,}$ (2)
[…]
$\bm{S}_{\text{sea}}(\varepsilon_{\text{F}})=\sum_{n}\int\frac{\text{d}^{2}\bm{k}}{(2\pi)^{2}}\,\mathit{f}_{n,\bm{k}}\delta\bm{s}_{n,\bm{k}}\text{ ,}$ (3)
[…]
Figure 1: (a) Band structure exhibiting spin-split Dirac cones at energies $\mp J_{\text{ex}}/2$ along the out-of-plane axis $\hat{\bm{z}}$, and a band gap at charge neutrality (yellow highlight), with $J_{\text{ex}}=40\,\text{meV}$ and $\lambda_{\text{KM}}=\lambda_{\text{R}}=4\,\text{meV}$ (b) Spin texture of the valence band, showing a sharp quenching at the band edge (yellow highlight) due to Kane-Mele SOC. (c) Damping-like non-equilibrium spin texture (arbitrary units) of the valence band, showing a source of DL torque at the band edge and at the Dirac point.
@@
We offer an explanation for the origin of the SOT plateau based on the semi-classical spin-pseudospin dynamics, relating it to band structure features. As shown in Fig. 1-(a), the band gap at charge neutrality (yellow highlight) is generated by the avoided crossing between the upper band of the spin majority Dirac cone and the lower band of the spin minority one, with positive and negative out-of-plane spin polarization respectively and both exhibiting the same helicity. For $\lambda_{\text{KM}}=0$ these bands present a radial pseudospin polarization along $\pm\hat{\bm{k}}$, opposite to one another, belonging to orthogonal pseudospin subspaces. Thus, the hybridization between both bands near the gap leads to a loss of their out-of-plane spin polarization due to spin-pseudospin entanglement, while their in-plane spin texture is unaffected. On the other hand, when introducing Kane-Mele SOC both bands acquire the same out-of-plane pseudospin component proportional to $\lambda_{\text{KM}}$, and thus share a common pseudospin subspace. In this case the hybridization between bands of up and down spin polarization stemming from the opposite Dirac cones modifies the in-plane spin texture [24]. This process results in a sharp variation of the spin helicity at the band edges, as highlighted in yellow in Fig. 1-(b), where the mark of strong spin-momentum locking forecasts a large $S_{\text{DL}}$. Indeed, as the Bloch states are transported under an applied electric field, those near the band edge feel a strong variation of the effective magnetic field, which drives the non-equilibrium spin dynamics resulting in a non-equilibrium spin texture source at the band edge, shown in Fig. 1-(c). This effect is opposite at the valence and conduction bands, manifesting broken particle-hole symmetry, and therefore, within the gap the fully filled valence band generates the DL SOT plateau [24].